Complete the Budgeted Income Statements below based on the information given - Check figures are given below6 7 Check figures are: January Budgeted Gross Margin = $62,650; February Budgeted 8 Operating Income = $23,442; March BudgeteА B C D E F G H I J K L M N O P Q R S T U V V 19 20 Other Operating Expenses 21 Commission Expense 22 Marketing Expense 23 In4 A relatively new company, Get Laughy Taffy, Inc., has determined they need to engage in a budgeting 5 process as part of thD E F G H I J K L M N O P A B C 35 Actual Purchases made in December 2019 $137,839 Purchases paid in current month 30.00% 3753 The annual interest rate on the long term debt is 4.00% Interest is incurred on the principal balance outstanding each monA E GH. M N UP R 3 4. Use the information provided on the December 31, 2019 Balance 5 Sheet shown below to help you create thG H I J K L M N O Q R S - B 450 7,665 14,977 131,549 А 33 Interest Payable 34 Marketing Payable 35 Income Taxes Payable 36 Di
Answer #1

Mentioned basis as per above in the particular for the values

Example

Cost of goods sold Jan 2020 - 179000*65% = 116,350/-

Commission Jan 2020 - 179,000*2.5%=4,475/-

Marketing expenses Jan 2020 - 3,540 + (179,000*1.5%) =6,225/-

Selling expenses Jan 2020 - 19800+(179,000*2%) = 23,380/-

Budged income statement of M/s Get Laughy Taffy Inc.
Particular Jan 31, 2020 Feb 29,2020 Mar 31,2020
Sales        1,79,000        1,84,000        1,63,000
Cost of sales (65% on sales)        1,16,350        1,19,600        1,05,950
Gross Margin            62,650            64,400            57,050
Other Operating expenses
Commission expenses (2.5% on sales)              4,475              4,600              4,075
Marketing expenses (3540+1.5%on sales)              6,225              6,300              5,985
insurance expenses (Jan 1305, Feb and March 1278 per month (15336/12)              1,305              1,278              1,278
Selling & Administrative expenses (19800+2%on sales)            23,380            23,480            23,060
Depreciation expenses              5,300              5,300              5,300
Total other Operating expenses            40,685            40,958            39,698
Operating income            21,965            23,442            17,352
Other income (Expenses)
Interest expenses (Since interest is calculated on prinicple O/s only. Accrued interest in not considered)
                 450                  450                  450
Income before income tax            21,515            22,992            16,902
Income tax Expenses              4,518              4,828              3,549
Net income            16,997            18,164            13,353
Dividend declared and paid in Feb            15,000
Transfer to reserves            16,997              3,164            13,353